### v1

# Gravitational lensing inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits

Abstract
It is believed that rays bent inside and outside photon spheres could affect partially the black hole shadow images by the Event Horizon Telescope and the rays near photon spheres would be detected by near-future space observations. The features of the rays near the photon spheres in not only black hole spacetimes but also exotic spacetimes would be important since one will need them to exclude black-hole mimickers. The deflection angles of the rays deflected by photon spheres diverge logarithmically and we can treat them by a strong-deflection-limit analysis. The error of the strong-deflection-limit analysis becomes large if antiphoton spheres exist in the spacetimes and the analysis breaks down when the photon spheres and the antiphoton spheres degenerate to form a marginally unstable photon sphere. This is because the deflection angles of the rays bent by the marginally unstable photon sphere diverge in powers. In this paper, we extend Eiroa, Romero, and Torres’s method to gravitational lensing of rays inside and outside of the marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits and we apply it to a Reissner-Nordström spacetime and a Hayward spacetime with the marginally unstable photon sphere. We have also confirmed that the deflection angles in the strong deflection limits by the method converge correctly to the deflection angle without approximations, while there are the mismatches of the coefficient of the power-divergent term of the deflection angles of the rays deflected just outside of the marginally unstable photon sphere in a semi-analytic calculation by the author previously.
I Introduction
Recently, the direct detection of gravitational waves by binary black hole has been reported by the LIGO and VIRGO Collaborations Abbott:2016blz ; LIGOScientific:2018mvr ; LIGOScientific:2020ibl ; KAGRA:2021vkt ; LIGOScientific:2025slb and the ring-shape images of supermassive black hole candidates at the centers of the giant elliptical galaxy M87 and the Milky way galaxy have been reported by the Event Horizon Telescope Collaboration Akiyama:2019cqa ; EventHorizonTelescope:2022wkp . Strong gravity near black holes would cause high energy phenomena and it would be important to detect the black holes and to understand their astrophysical environments.
Compact objects such as black holes and wormholes have photon spheres which are spheres constructed by unstable circular light orbits due to their strong gravity Perlick_2004_Living_Rev . Several aspects of the photon spheres such as the bounds of their radii Hod:2017xkz ; Peng:2018nkj ; Hod:2020pim ; Hod:2023jmx , their numbers Hod:2017zpi ; Cunha:2020azh ; Cunha:2024ajc ; Padhye:2024yrj , and their relationship to photon absorption cross sections Sanchez:1977si ; Decanini:2010fz ; Wei:2011zw , quasinormal modes Press:1971wr ; Goebel_1972 ; Stefanov:2010xz ; Raffaelli:2014ola ; Igata:2025plb ; Igata:2025hpy , a centrifugal force and gyroscopic precession Abramowicz_Prasanna_1990 ; Abramowicz:1990cb ; Allen:1990ci ; Hasse_Perlick_2002 , Bondi’s sonic horizon Mach:2013gia ; Chaverra:2015bya ; Cvetic:2016bxi ; Koga:2016jjq ; Koga:2018ybs ; Koga:2019teu , and the stability of thin-shell wormholes Barcelo:2000ta ; Koga:2020gqd ; Tsukamoto:2023kvk , were investigated eagerly. If the gravity of compact objects is moderate, they could have also antiphoton spheres which are spheres formed by stable circular light orbits, but the instability of the compact objects may be caused by the slow decay of linear waves near the antiphoton spheres Keir:2014oka ; Cardoso:2014sna ; Cunha:2017qtt ; Cunha:2017eoe ; Cunha:2022gde ; Zhong:2022jke . The alternatives to the photon spheres and the antiphoton spheres were also suggested by several researchers Claudel:2000yi ; Koga:2019uqd ; Cunha:2017eoe ; Gibbons:2016isj ; Shiromizu:2017ego ; Yoshino:2017gqv ; Galtsov:2019bty ; Galtsov:2019fzq ; Siino:2019vxh ; Yoshino:2019dty ; Cao:2019vlu ; Yoshino:2019mqw ; Lee:2020pre .
Gravitational lenses are often used to survey dark objects such as the black holes, extrasolar planets, and dark matters Schneider_Ehlers_Falco_1992 ; Schneider_Kochanek_Wambsganss_2006 . The gravitational lenses have been mainly studied in a weak gravitational field for a hundred years since we do not have enough technology to detect astrophysical phenomena in strong gravitational fields until the last decade. Rays deflected by the photon spheres of the black holes and their gravitational lensing effects in strong gravitational fields have been intermittently investigated Hagihara_1931 ; Darwin_1959 ; Atkinson_1965 ; Luminet_1979 ; Ohanian_1987 ; Nemiroff_1993 ; Frittelli_Kling_Newman_2000 ; Virbhadra_Ellis_2000 ; Bozza:2001xd ; Bozza:2002zj ; Perlick:2003vg ; Nandi:2006ds ; Virbhadra:2008ws ; Hioki:2009na ; Bozza_2010 ; Tsupko:2017rdo ; Nakao:2018knn ; Gralla:2019xty ; Okabayashi:2020apz ; Hsieh:2021scb ; Aratore:2021usi ; Tsupko:2022kwi ; Aratore:2024bro ; Tsupko:2025hhf . The effects of the photon spheres on the apparent shape of stars collapsing into the black holes also have been investigated Ames_1968 ; Synge:1966okc ; Yoshino:2019qsh ; Koga:2025dii . It is believed that the rays deflected just inside and outside photon spheres could affect partially the observed ring images of the supermassive black hole candidates at the center of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp and the rays bent near the photon spheres would be detected by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr . It is important to investigate the features of the rays near the photon spheres in not only black hole spacetimes but also exotic compact objects such as wormholes Ellis:1973yv ; Chetouani:1984qdm ; Perlick:2003vg ; Muller:2004dq ; Nandi:2006ds ; Muller:2008zza ; Tsukamoto:2012xs ; Ohgami:2015nra ; Tsukamoto:2016qro ; Tsukamoto:2016zdu ; Shaikh:2018kfv ; Shaikh:2018oul ; Bronnikov:2018nub ; Shaikh:2019jfr ; Bronnikov:2021liv ; Maeda:2021wnl ; Olmo:2023lil ; Bugaev:2023mlc ; Zhang:2024sgs ; Xavier:2024iwr ; Tsukamoto:2024pid ; Bugaev:2024dte ; Solodukhin:2025opw and naked singularity Virbhadra:1998dy ; Virbhadra:2002ju ; Hioki:2009na ; Shaikh:2018lcc ; Shaikh:2019itn since one needs to exclude exotic compact objects or black-hole mimickers to reinforce claims that the observed objects must be supermassive black holes.
Gravitational lensing of rays reflected by just outside of a photon sphere in a Schwarzschild spacetime in a strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$, where $b$ and $b_{\mathrm{m}}$ an impact parameter and a critical impact parameter of the ray, by Bozza et al. Bozza:2001xd . The deflection angle of the rays which pass near outside of the photon sphere diverges logarithmically can be expressed by
$\alpha$ $=$ $-\bar{a}_{+}\log\left(\frac{b}{b_{\mathrm{m}}}-1\right)+\bar{b}_{+},$ (1)
where $\bar{a}_{+}$ and $\bar{b}_{+}$ are a coefficient and a term which are determined by the metric of the spacetime with the photon sphere. Eiroa, Romero, and Torres suggested a method to obtain the coefficient $\bar{a}_{+}$ and the term $\bar{b}_{+}$ numerically and they applied it to a Reissner-Nordström black hole spacetime Eiroa:2002mk while Bozza made a semi-analytic formula in a general, static, spherically symmetric, and asymptotically-flat spacetime with the photon sphere to obtain the coefficient $\bar{a}_{+}$ and the term $\bar{b}_{+}$ and got $\bar{a}_{+}$ in an analytic form and $\bar{b}_{+}$ partially numerically in the Reissner-Nordström black hole spacetime Bozza:2002zj . The strong-deflection-limits analysis has been intensively studied and it has been extended over the last two decades  Bozza:2002af ; Eiroa:2002mk ; Petters:2002fa ; Eiroa:2003jf ; Bozza:2004kq ; Bozza:2005tg ; Bozza:2006sn ; Bozza:2006nm ; Iyer:2006cn ; Bozza:2007gt ; Tsukamoto:2016qro ; Ishihara:2016sfv ; Tsukamoto:2016oca ; Tsukamoto:2016zdu ; Tsukamoto:2016jzh ; Shaikh:2018oul ; Shaikh:2019itn ; Takizawa:2021gdp ; Zhang:2024sgs ; Igata:2025taz . Tsukamoto Tsukamoto:2016oca ; Tsukamoto:2016jzh investigated an alternative of Bozza’s method Bozza:2002zj , obtained $\bar{a}_{+}$ and $\bar{b}_{+}$ in analytic forms in the Reissner-Nordström spacetime and confirmed that they are consistent with $\bar{a}_{+}$ and $\bar{b}_{+}$ obtained by Eiroa, Romero, and Torres Eiroa:2002mk and by Bozza Bozza:2002zj .
If we consider gravitational lensing effects of rays passing inside and outside of the photon spheres in spacetimes without event horizons Shaikh:2019itn ; Shaikh:2019jfr , the deflection angles of the rays in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ can be expressed by
$\alpha$ $=$ $-\bar{a}_{\pm}\log\left|\frac{b}{b_{\mathrm{m}}}-1\right|+\bar{b}_{\pm},$ (2)
where $\bar{a}_{-}$ and $\bar{b}_{-}$ are a coefficient and a term, respectively, for the ray reflected near inside of the photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}-0$ and they can be obtained from the metric of the spacetimes. Shaikh et al. Shaikh:2019itn obtained $\bar{a}_{-}$ as an analytical form and $\bar{b}_{-}$ numerically in a Reissner-Nordström naked singularity spacetime and the analytical form of $\bar{b}_{-}$ was obtained by Tsukamoto Tsukamoto:2021fsz .
The errors of deflection angles in the strong deflection limits becomes large if antiphoton spheres exist in the spacetimes and the strong-deflection-limit analysis for the photon spheres breaks down when the photon spheres and the antiphoton spheres degenerate to be a marginally unstable photon sphere since the deflection angles of the rays bent by the marginally unstable photon sphere in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ diverge in powers. In this paper, we treat the deflection angle of the rays in the following form:
$\alpha(b)=\frac{\bar{c}_{\pm}}{\left|\frac{b}{b_{\mathrm{m}}}-1\right|^{\frac{1}{6}}}+\bar{d}_{\pm},$ (3)
where $\bar{c}_{+}$ and $\bar{d}_{+}$ ($\bar{c}_{-}$ and $\bar{d}_{-}$) are a coefficient and a term, respectively, for the rays reflected by just outside (inside) of the marginally unstable photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$(0-0). In Ref. Tsukamoto:2020iez , Tsukamoto extended Bozza’s method to gravitational lensing of the rays reflected by just outside of the marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ and applied it to the Reissner-Nordström naked singularity spacetime with the marginally unstable photon sphere. Recently, Sasaki Sasaki:2025web investigated the details of the gravitational lensing of the rays reflected inside and outside of the marginally unstable photon sphere in the Reissner-Nordström naked singularity spacetime and obtained the higher order terms of deflection angles in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$. Sasaki Sasaki:2025web also pointed out that the deflection angle by Tsukamoto Tsukamoto:2020iez in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ does not converge to the deflection angle with no approximations due to the invalid value of the coefficient $\bar{c}_{+}$.
In this paper, we extend Eiroa, Romero, and Torres’s method Eiroa:2002mk to gravitational lensing of rays just inside and outside of the marginally unstable photon sphere in the general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and we apply it to the Reissner-Nordström spacetime and a Hayward spacetime Hayward:2005gi with the marginally unstable photon sphere. We also confirm that mismatches in the coefficients $\bar{c}_{+}$ which were calculated by the author in the semi-analytic method previously Tsukamoto:2020iez , are modified in the new approach.
This paper is organized as follows. We review the deflection angle of rays in Sec. II, we investigate extended Eiroa, Romero, and Torres’s method to obtain the deflection angle of the rays inside and outside of the marginally unstable photon sphere in the strong deflection limits in the general, static, spherically symmetric, and asymptotically-flat spacetime and we apply it to the Reissner-Nordström spacetime and the Hayward spacetime with the marginally unstable photon spheres in Sec. III, and we investigate observable in gravitational lensing by using the deflection angles and a lens equation in Sec. IV. We conclude our results in Sec. V. We use a unit that a light speed and Newton’s constant are unity in this paper.
II Deflection angle of rays
We consider a general, asymptotically flat, static and spherically symmetric spacetime with a line element, in coordinates $\infty<t<+\infty$, $r<+\infty$, $0\leq\vartheta\leq\pi$, and $0\leq\varphi<2\pi$,
$\mathrm{d}s^{2}=-A(r)\mathrm{d}t^{2}+B(r)\mathrm{d}r^{2}+C(r)(\mathrm{d}\vartheta^{2}+\sin^{2}\vartheta\mathrm{d}\varphi^{2}),$ (4)
with asymptotically flat conditions
$\lim_{r\rightarrow\infty}A(r)=\lim_{r\rightarrow\infty}B(r)=1+O(r^{-1}),$ (5)
$\lim_{r\rightarrow\infty}C(r)=r^{2}+O(r)$ (6)
and we assume that $A(r)$, $B(r)$, and $C(r)$ are positive and finite in a region where we consider. We also assume that the spacetime has circular light orbits at $r=r_{\mathrm{m}}$, or a condition
$D_{\mathrm{m}}=\frac{C^{\prime}_{\mathrm{m}}}{C_{\mathrm{m}}}-\frac{A^{\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}=0$ (7)
holds there. Here, we have defined $D(r)$ as
$D(r)\equiv\frac{C^{\prime}(r)}{C(r)}-\frac{A^{\prime}(r)}{A(r)},$ (8)
where denotes a differentiation with respect to $r$ and functions with the subtract m denotes the functions at $r=r_{\mathrm{m}}$. The circular light orbits are unstable (stable) if a condition
$D_{\mathrm{m}}^{\prime}=\frac{C_{\mathrm{m}}^{\prime\prime}}{C_{\mathrm{m}}}-\frac{A^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}>0\quad(<0)$ (9)
is held and we name a sphere formed by the unstable (stable) circular light orbits a photon (an antiphoton) sphere. We note that the derivatives of $D(r)$ are give by
$D^{\prime}(r)$ $=$ $\frac{C^{\prime\prime}(r)}{C(r)}-\frac{C^{\prime}(r)^{2}}{C(r)^{2}}$ (10)
$-\frac{A^{\prime\prime}(r)}{A(r)}+\frac{A^{\prime}(r)^{2}}{A(r)^{2}}$
and
$D^{\prime\prime}(r)$ $=$ $\frac{C^{\prime\prime\prime}(r)}{C(r)}-3\frac{C^{\prime\prime}(r)C^{\prime}(r)}{C(r)^{2}}+2\frac{C^{\prime}(r)^{3}}{C(r)^{3}}$ (11)
$-\frac{A^{\prime\prime\prime}(r)}{A(r)}+3\frac{A^{\prime\prime}(r)A^{\prime}(r)}{A(r)^{2}}-2\frac{A^{\prime}(r)^{3}}{A(r)^{3}}.$
The spacetime has time-translational and axial Killing vectors $t^{\mu}\partial_{\mu}=\partial_{t}$ and $\varphi^{\mu}\partial_{\mu}=\partial_{\varphi}$ due to stationarity and axial symmetry of the spacetime, respectively, and we can assume $\vartheta=\pi/2$ without loss of generality because of the spherical symmetry of the spacetime.
From the wave number of a ray $k^{\mu}\equiv\dot{x}^{\mu}$, where $x^{\mu}$ is the coordinates and the dot denotes a differentiation with respect to an affine parameter along the trajectory of the ray, the trajectory of the ray is obtained as
$A(r)\dot{t}^{2}+B(r)\dot{r}^{2}+C(r)\dot{\varphi}^{2}=0.$ (12)
We consider that the ray, which comes from a spatial infinity, is reflected by a lensing object at the closest distance $r=r_{0}$ and is away to the spatial infinity. In this paper, we do not consider a case that the rays coming from the spatial infinity, passes through a wormhole throat, and leave to another spatial infinity if we consider wormhole spacetimes. At the closest distance $r=r_{0}$, due to
$\dot{r}_{0}\equiv\left.\dot{r}\right|_{r=r_{0}}=0,$ (13)
where functions with the subtract 0 denote the functions at the closest distance $r=r_{0}$, we obtain, from Eq. (12),
$A_{0}\dot{t}_{0}^{2}=C_{0}\dot{\varphi}^{2}_{0}.$ (14)
From Eq. (14), the impact parameter $b$ of the ray defined as $b\equiv L/E$, where $E\equiv-g_{\mu\nu}t^{\mu}k^{\nu}$ and $L\equiv g_{\mu\nu}\varphi^{\mu}k^{\nu}$ are the conserved energy and the angular momentum of the ray, respectively, is expressed as
$b=b(r_{0})=\frac{L}{E}=\frac{C_{0}\dot{\varphi}_{0}}{A_{0}\dot{t}_{0}}=\sqrt{\frac{C_{0}}{A_{0}}},$ (15)
and it is constant along the trajectory of the ray. For simplicity, unless otherwise stated, we assume that the value of $b$ is positive.
By rescaling the affine parameter of the ray, the trajectory (12) is rewritten in
$\dot{r}^{2}+V(r,b)=0,$ (16)
where $V(r,b)$ is an effective potential for the radial motion of the ray defined by
$V(r,b)\equiv\frac{1}{B(r)}\left(\frac{b^{2}}{C(r)}-\frac{1}{A(r)}\right).$ (17)
The ray can be in regions, where the effective potential is nonnegative, i.e., $V(r,b)\leq 0$. From Eqs. (5) and (6), we obtain $\lim_{r\rightarrow\infty}V(r)=-1<0$. Thus, the ray can be at the spatial infinity.
The trajectory of the ray (12) can be rewritten in
$\left(\frac{\mathrm{d}r}{\mathrm{d}\varphi}\right)^{2}=\frac{C(r)}{B(r)}\left(\frac{C(r)}{A(r)b^{2}}-1\right)=-\frac{C^{2}(r)}{B(r)b^{2}}V(r,b)$ (18)
and we obtain the deflection angle $\alpha(r_{0})$ of the ray as
$\alpha(r_{0})$ $\equiv$ $2\int^{\infty}_{r_{0}}\frac{\mathrm{d}r}{\sqrt{\frac{C(r)}{B(r)}\left(\frac{C(r)}{A(r)b^{2}(r_{0})}-1\right)}}-\pi$ (19)
$=$ $2\int^{\infty}_{r_{0}}\frac{b(r_{0})\sqrt{B(r)}\mathrm{d}r}{C(r)\sqrt{-V(r,b(r_{0}))}}-\pi.$
III Deflection angle just inside and outside of the marginally unstable photon sphere in the strong deflection limits
We extend Eiroa, Romero, and Torres’s approach Eiroa:2002mk to the marginally unstable photon sphere in the strong deflection limits $r_{0}\rightarrow r_{\mathrm{m}}\pm 0$ or $b(r_{0})\rightarrow b_{\mathrm{m}}\pm 0$, where $b_{\mathrm{m}}\equiv b(r_{\mathrm{m}})$ is the critical impact parameter of the ray. Here and hereinafter, the upper (lower) sign is correspond with the image of rays reflected by outside (inside) of the marginally unstable photon sphere.
We assume that conditions
$D_{\mathrm{m}}=\frac{C^{\prime}_{\mathrm{m}}}{C_{\mathrm{m}}}-\frac{A^{\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}=0,$ (20)
$D^{\prime}_{\mathrm{m}}=\frac{C^{\prime\prime}_{\mathrm{m}}}{C_{\mathrm{m}}}-\frac{A^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}=0,$ (21)
and
$D^{\prime\prime}_{\mathrm{m}}=\frac{C^{\prime\prime\prime}_{\mathrm{m}}}{C_{\mathrm{m}}}-\frac{A^{\prime\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}>0$ (22)
is satisfied so that the spacetime has the marginally unstable photon sphere. Under the assumptions, we obtain
$V(r_{\mathrm{m}},b_{\mathrm{m}})=V^{\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=V^{\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=0,$ (23)
$V^{\prime\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=-\frac{b_{\mathrm{m}}D^{\prime\prime}_{\mathrm{m}}}{B_{\mathrm{m}}C_{\mathrm{m}}}=-\frac{D^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}B_{\mathrm{m}}}<0,$ (24)
and the deflection angle of the ray bent outside and inside of the marginally unstable photon sphere in the strong deflection limits $r_{0}\rightarrow r_{\mathrm{m}}\pm 0$ or $b\rightarrow b_{\mathrm{m}}\pm 0$ will be expressed by,
$\alpha=\frac{c_{\pm}}{\left|r_{0}-r_{\mathrm{m}}\right|^{1/2}}+\bar{d}_{\pm},$ (25)
where $c_{\pm}$ is the coefficient and $\bar{d}_{\pm}$ is the term. From the relation
$\lim_{r_{0}\rightarrow r_{\mathrm{m}}\pm 0}\left[\alpha-\frac{c_{\pm}}{\left|r_{0}-r_{\mathrm{m}}\right|^{1/2}}-\bar{d}_{\pm}\right]=0,$ (26)
where $\alpha$ is given by Eq. (19), and its derivative with respective to $r_{0}$ given by
$\lim_{r_{0}\rightarrow r_{\mathrm{m}}\pm 0}\left[\frac{\mathrm{d}\alpha}{\mathrm{d}r_{0}}\pm\frac{c_{\pm}}{2\left|r_{0}-r_{\mathrm{m}}\right|^{3/2}}\right]=0,$ (27)
we can obtain the coefficient $c_{\pm}$ and the term $\bar{d}_{\pm}$ as, in numerical,
$c_{\pm}=\lim_{r_{0}\rightarrow r_{\mathrm{m}}\pm 0}\left[\mp 2\left|r_{0}-r_{\mathrm{m}}\right|^{3/2}\frac{\mathrm{d}\alpha}{\mathrm{d}r_{0}}\right]$ (28)
and
$\bar{d}_{\pm}$ $=$ $\lim_{r_{0}\rightarrow r_{\mathrm{m}}\pm 0}\left[\alpha-\frac{c_{\pm}}{\left|r_{0}-r_{\mathrm{m}}\right|^{1/2}}\right]$ (29)
$=$ $\lim_{r_{0}\rightarrow r_{\mathrm{m}}\pm 0}\left(\alpha\pm 2\left|r_{0}-r_{\mathrm{m}}\right|\frac{\mathrm{d}\alpha}{\mathrm{d}r_{0}}\right),$
respectively.
By using series of the expansion of $b(r_{0})$ around $r_{0}=r_{\mathrm{m}}$,
$b(r_{0})=b_{\mathrm{m}}+\frac{b_{\mathrm{m}}D^{\prime\prime}_{\mathrm{m}}}{12}\left(r_{0}-r_{\mathrm{m}}\right)^{3}+O\left(\left(r_{0}-r_{\mathrm{m}}\right)^{4}\right),$ (30)
the deflection angle (25) can be expressed by
$\alpha(b)=\frac{\bar{c}_{\pm}}{\left|\frac{b}{b_{\mathrm{m}}}-1\right|^{1/6}}+\bar{d}_{\pm},$ (31)
where $\bar{c}_{\pm}$ is given by
$\bar{c}_{\pm}\equiv\left|\frac{D^{\prime\prime}_{\mathrm{m}}}{12}\right|^{1/6}c_{\pm}.$ (32)
III.1 Reissner-Nordström spacetime
In the Reissner-Nordström black hole spacetime, gravitational lensing of rays bent by outside of a photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ was investigated by Eiroa, Romero, and Torres Eiroa:2002mk in a numerical way and by Bozza Bozza:2002zj in a semi-analytical way, and extended source effects Eiroa:2002mk , the time delay of rays Sereno:2003nd , retrolensing Eiroa:2003jf ; Tsukamoto:2016oca ; Tsukamoto:2016jzh , and higher order terms of the deflection angle Tsukamoto:2022tmm ; Tsukamoto:2022uoz ; Sasaki:2023rdf were studied. The images of light sources at near the photon sphere in the Reissner-Nordström black hole spacetime were investigated in Ref. Aratore:2024bro ; Tsupko:2025hhf by using a strong-deflection-limit analysis with arbitrary source distances Bozza:2007gt . In the Reissner-Nordström naked singularity spacetime, the size of the photon sphere or the shadows Zakharov:2014lqa and gravitational lensing of rays deflected inside and outside of a photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}\pm 0$ were investigated in Refs. Shaikh:2019itn ; Tsukamoto:2021fsz ; Tsukamoto:2021lpm .
Tsukamoto Tsukamoto:2020iez got the deflection angle with the coefficient
$\bar{c}_{+}=2^{5/3}3^{1/2}\sim 5.49892$ (33)
and the constant
$\bar{d}_{+}=-\sqrt{6}-\pi\sim-5.59108$ (34)
in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ in the semi-analytical method. Sasaki Sasaki:2025web investigated the deflection angle of the rays reflected inside and outside of the marginally unstable photon sphere the Reissner-Nordström naked singularity spacetime in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and obtained
$\bar{c}_{+}=\frac{\Gamma\left(\frac{1}{12}\right)\Gamma\left(\frac{7}{12}\right)}{2^{1/6}\sqrt{3}\Gamma\left(\frac{2}{3}\right)}\sim 6.67748,$ (35)
$\bar{d}_{+}=-\sqrt{6}-\pi\sim-5.59108,$ (36)
$\bar{c}_{-}=\sqrt{3}\bar{c}_{+}\sim 11.5657,$ (37)
and
$\bar{d}_{-}=\bar{d}_{+}=-\sqrt{6}-\pi\sim-5.59108$ (38)
where $\Gamma\left(z\right)$ is the gamma function, and higher terms in the deflection angles. Sasaki pointed out that $\bar{c}_{+}$ in Eq. (33) is invalid since that the deflection angle with $\bar{c}_{+}$ in Eq. (33) and $\bar{d}_{+}$ in Eq. (34) obtained by Tsukamoto Tsukamoto:2020iez does not converge to the deflection angle without approximations.
The Reissner-Nordström spacetime with the functions of the metric
$A(r)=B(r)^{-1}=1-\frac{2M}{r}+\frac{9M^{2}}{8r^{2}},$ (39)
$C(r)=r^{2},$ (40)
has the marginally unstable photon sphere, which is formed by rays with the critical impact parameter
$b_{\mathrm{m}}=\frac{3\sqrt{6}M}{2},$ (41)
at $r=r_{\mathrm{m}}$, where $r_{\mathrm{m}}$ is given by
$r_{\mathrm{m}}=\frac{3M}{2},$ (42)
and it holds the conditions
$D_{\mathrm{m}}=D^{\prime}_{\mathrm{m}}=0,$ (43)
$D^{\prime\prime}_{\mathrm{m}}=\frac{64}{9}>0.$ (44)
Figure 1 shows the effective potential $V(r/M,b)$ for the radial motion of the rays.
Figure 1: The effective potential $V(r/M,b)$ in the Reissner-Nordström spacetime with the marginally unstable photon sphere is shown. The broken (green), solid (red), and dotted (magenta) curves denote $V(r/M,b)$ for $b=0.99b_{\mathrm{m}}$, $b_{\mathrm{m}}$, and $1.01b_{\mathrm{m}}$, respectively.
From Eqs. (28), (29), and (32), we obtain
$\bar{c}_{+}\sim 6.67748$ (45)
and
$\bar{d}_{+}\sim-5.59011$ (46)
numerically for the rays just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime.
From Eqs. (28) and (29), for the rays just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime, we get
$\bar{c}_{-}\sim 11.5658$ (47)
and
$\bar{d}_{-}\sim-5.64071$ (48)
in numerical.
The deflection angle $\alpha$ is shown in Fig. 2 and the percent error of the deflection angle of Eq. (3.12) defined by
$\frac{\alpha\>\mathrm{of\>Eq.}(3.12)-\alpha\>\mathrm{of\>Eq.}(2.17)}{\alpha\>\mathrm{of\>Eq.}(2.17)}\times 100$ (49)
is shown in Fig. 3.
Figure 2: The deflection angles $\alpha$ in the Reissner-Nordström spacetime by using Eq. (2.17) is denoted by the dot-dashed (magenta) curve. The one by using Eq. (3.12) with $\bar{c}_{+}$ of Eq. (45), $\bar{d}_{+}$ of Eq. (46), $\bar{c}_{-}$ of Eq. (47), and $\bar{d}_{-}$ of Eq. (48) in our numerical method, with $\bar{c}_{+}$ of Eq. (35), $\bar{d}_{+}$ of Eq. (36), $\bar{c}_{-}$ of Eq. (37), and $\bar{d}_{-}$ of Eq. (38) by Sasaki Sasaki:2025web , and with $\bar{c}_{+}$ of Eq. (33) and $\bar{d}_{+}$ of Eq. (34) by Tsukamoto Tsukamoto:2020iez are denoted by the dashed (green), dotted (black), and solid (red) curves, respectively. Notice that the dashed (green) and dotted (black) curves are overlapped.
Figure 3: The percent errors of deflection angles $\alpha$ of Eq. (3.12) against the deflection angle $\alpha$ of Eq. (2.17) in the Reissner-Nordström spacetime are plotted in top and bottom panels for rays producing images outside and inside of the marginally unstable photon sphere, respectively. In the top pnnel, dashed (green), dotted (black), solid (red) curves denote the percent errors of deflection angles $\alpha$ of Eq. (3.12) with $\bar{c}_{+}$ of Eq. (45) and $\bar{d}_{+}$ of Eq. (46) in our numerical method, with $\bar{c}_{+}$ of Eq. (35) and $\bar{d}_{+}$ of Eq. (36) by Sasaki Sasaki:2025web , and with $\bar{c}_{+}$ of Eq. (33) and $\bar{d}_{+}$ of Eq. (34) by Tsukamoto Tsukamoto:2020iez , respectively. In the bottom panel, dashed (green) and dotted (black) curves denote the percent errors of deflection angles $\alpha$ of Eq. (3.12) with $\bar{c}_{-}$ of Eq. (47) and $\bar{d}_{-}$ of Eq. (48) in our numerical method, and with $\bar{c}_{-}$ of Eq. (37) and $\bar{d}_{-}$ of Eq. (38) by Sasaki Sasaki:2025web , respectively. Note that the dashed (green) and dotted (black) curves are overlapped.
III.2 Hayward spacetime
The shadow images of the photon sphere of Hayward black holes Li:2013jra ; Tsukamoto:2017fxq and the deflection angle of a light in a weak-field approximation and in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ in the Hayward black hole spacetime Wei:2015qca were studied. Chiba and Kimura Chiba:2017nml showed the shadow images in black hole and over-charged cases.
Chiba and Kimura Chiba:2017nml investigated the deflection angle of the rays reflected by the marginally unstable photon sphere and obtained the coefficient $\bar{c}_{+}$ as, from Eqs. (21) and (22) in Ref. Chiba:2017nml ,
$\bar{c}_{+}=\frac{c_{+}M^{1/6}}{b_{\mathrm{m}}^{1/6}}\sim 6.01316,$ (50)
where
$c_{+}=2^{11/6}3^{2/3}5^{5/4}\int^{\infty}_{0}\frac{dy}{\sqrt{432y+900y^{2}+625y^{3}}}\sim 7.771.$ (51)
but they do not consider the constants $\bar{d}_{+}$ and $\bar{d}_{-}$ and the coefficient $\bar{c}_{-}$. Tsukamoto Tsukamoto:2020iez studied the deflection angle of the rays reflected outside of a marginally unstable photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ and got the coefficient as, by using the semi-analytic method,
$\bar{c}_{+}\sim\frac{2^{13/6}3^{1/3}}{5^{1/6}}\sim 4.95196$ (52)
and the constant
$\bar{d}_{+}=-4\sqrt{\frac{6}{5}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi\sim-5.62607.$ (53)
The percent errors of the deflection angle of Eq. (3.12) with the coefficients $\bar{c}_{+}$ and the constants $\bar{d}_{+}$ are plotted in Fig. 6.
The Hayward spacetime Hayward:2005gi has the marginally unstable photon sphere at $r=r_{\mathrm{m}}=25M/12$ made by the rays with the critical impact parameter
$b_{\mathrm{m}}=\frac{25\sqrt{5}M}{12}$ (54)
and it holds the conditions
$D_{\mathrm{m}}=D^{\prime}_{\mathrm{m}}=0,$ (55)
$D^{\prime\prime}_{\mathrm{m}}=\frac{1728}{625}>0$ (56)
when the functions of the metric are
$A(r)=B(r)^{-1}=1-\frac{2Mr^{2}}{r^{3}+2q^{2}M},$ (57)
$C(r)=r^{2},$ (58)
where we have set $2q^{2}M=5^{5}M^{3}/(2^{6}3^{3})$. The effective potentials $V(r/M,b)$ of the rays for the radial motion are plotted in Fig. 4.
Figure 4: The effective potentials $V(r/M,b)$ for $b=0.99b_{\mathrm{m}}$, $b_{\mathrm{m}}$, and $1.01b_{\mathrm{m}}$ in the Hayward spacetime with the marginally unstable photon sphere are shown as the broken (green), solid (red), and dotted (magenta) curves, respectively.
From Eqs. (28), (29), and (32), we obtain
$\bar{c}_{+}\sim 6.01322$ (59)
and
$\bar{d}_{+}\sim-5.60584$ (60)
in numerical for the Hayward spacetime with the marginally unstable photon sphere. From Eqs. (28) and (29), we get the numerical values of
$\bar{c}_{-}\sim 10.4155$ (61)
and
$\bar{d}_{-}\sim-5.69675.$ (62)
The deflection angle $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere is shown in Fig. 5 and the percent error of the deflection angle of Eq. (3.12) with the coefficients $\bar{c}_{+}$ and the constants $\bar{d}_{+}$ is shown in Fig. 6.
Figure 5: The deflection angles $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere are shown. The dot-dashed (magenta) curve denotes the deflection angle $\alpha$ of Eq. (2.17). The dashed (green), dotted (black), and solid (red) curves denote the deflection angles $\alpha$ of Eq. (3.12) with $\bar{c}_{+}$ of Eq. (59), $\bar{d}_{+}$ of Eq. (60), $\bar{c}_{-}$ of Eq. (61), and $\bar{d}_{-}$ of Eq. (62) in our numerical method, with $\bar{c}_{+}$ of Eq. (50) and $\bar{d}_{+}=0$ by Chiba and Kimura Chiba:2017nml , and with $\bar{c}_{+}$ of Eq. (52) and $\bar{d}_{+}$ of Eq. (53) by Tsukamoto Tsukamoto:2020iez , respectively.
Figure 6: The percent errors of the deflection angles $\alpha$ of Eq. (3.12) against the deflection angle $\alpha$ of Eq. (2.17) in the Hayward spacetime are plotted in top and bottom panels for the rays producing images outside and inside of the marginally unstable photon sphere, respectively. In the top pannel, the dashed (green), dotted (black), and solid (red) curves denote the percent errors of deflection angles $\alpha$ of Eq. (3.12) with $\bar{c}_{+}$ of Eq. (59) and $\bar{d}_{+}$ of Eq. (60) in our numerical method, with $\bar{c}_{+}$ of Eq. (50) and $\bar{d}_{+}=0$ by Chiba and Kimura Chiba:2017nml , and with $\bar{c}_{+}$ of Eq. (52) and $\bar{d}_{+}$ of Eq. (53) by Tsukamoto Tsukamoto:2020iez , respectively. In the bottom panel, the dashed (green) curve denote the percent error of the deflection angle $\alpha$ of Eq. (3.12) with $\bar{c}_{-}$ of Eq. (61), and $\bar{d}_{-}$ of Eq. (62) in our numerical method.
We have confirmed the coefficient $\bar{c}_{+}$ (50) by Chiba and Kimura Chiba:2017nml and the constant $\bar{d}_{+}$ (53) by Tsukamoto Tsukamoto:2020iez are correct while the coefficient $\bar{c}_{+}$ (52) by using the semi-analytic method in Ref.Tsukamoto:2020iez should be modified.
IV Observable of gravitational lensing in the strong deflection limits
In this section, we consider the observable of gravitational lensing of the rays bent inside and outside of the marginally unstable photon sphere in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ or $r\rightarrow r_{\mathrm{m}}\pm 0$. We consider that a light source S with a source angle $\phi$ emits the ray with an impact parameter $b$, a lens L reflects the ray by a deflection angle $\alpha$ and an observer O sees the source S as an image I with an image angle $\theta$. The lens configuration is shown as Fig. 7.
Figure 7: A small lens configuration with an effective deflection angle $\left|\bar{\alpha}\right|\ll 1$, an image angle $\left|\theta\right|\ll 1$, and a source angle $\left|\phi\right|\ll 1$. A ray emitted by a light source S with an impact parameter $b$ is bent by a lens object L and it reaches to an observer O who observes an image I. $d_{\mathrm{OS}}$, $d_{\mathrm{LS}}$, and $d_{\mathrm{OL}}$ are angular distances between O and S, between L and S, and between O and L, respectively.
We introduce an effective deflection angle $\bar{\alpha}$ of the ray defined by
$\bar{\alpha}\equiv\alpha\quad\mathrm{mod}\;2\pi$ (63)
and the deflection angle $\alpha$ can be expressed by
$\alpha=\bar{\alpha}+2\pi n,$ (64)
where a non-negative integer $n$ is the winding number of the ray.
We assume that the angles $\left|\bar{\alpha}\right|$, $\left|\theta\right|$, and $\left|\phi\right|$ are small, i.e., $\left|\bar{\alpha}\right|,\left|\theta\right|,\left|\phi\right|\ll 1$. Then, a lens equation Bozza:2008ev is obtained as
$\theta=\frac{d_{\mathrm{LS}}}{d_{\mathrm{OS}}}\bar{\alpha}+\phi,$ (65)
where $d_{\mathrm{LS}}$ and $d_{\mathrm{OS}}$ are angular distances between L and S and between O and S, respectively, and the image angle $\theta$ can be expressed by
$\theta=\frac{b}{d_{\mathrm{OL}}},$ (66)
where $d_{\mathrm{OL}}=d_{\mathrm{OS}}-d_{\mathrm{LS}}$ is an angular distance between O and L.
We express the deflection angle $\alpha$ of the ray reflected just inside and outside of the marginally unstable photon sphere and its derivative respective to the image angle $\theta$ as
$\alpha(\theta)=\frac{\bar{c}_{\pm}}{\left|\frac{\theta}{\theta_{\infty}}-1\right|^{1/6}}+\bar{d}_{\pm},$ (67)
and
$\frac{d\alpha}{d\theta}=\mp\frac{\bar{c}_{\pm}}{6\theta_{\infty}\left|\frac{\theta}{\theta_{\infty}}-1\right|^{7/6}},$ (68)
respectively, where $\theta_{\infty}$ defined by
$\theta_{\infty}\equiv\frac{b_{\mathrm{m}}}{d_{\mathrm{OL}}}$ (69)
is an image angle for the marginally unstable photon sphere. By substituting $\theta^{0}_{n\pm}$, which is defined by
$\alpha\left(\theta^{0}_{n\pm}\right)=2\pi n$ (70)
for each positive winding number $n\geq 1$, into the deflection angle $\alpha(\theta)$ expressed by Eq. (67) of the ray reflected just inside and outside of the marginally unstable photon sphere, we obtain $\theta^{0}_{n\pm}$ as
$\theta^{0}_{n\pm}=\left[1\pm\left(\frac{\bar{c}_{\pm}}{2\pi n-\bar{d}_{\pm}}\right)^{6}\right]\theta_{\infty}.$ (71)
We expand the deflection angle $\alpha\left(\theta\right)$ around $\theta=\theta^{0}_{n\pm}$ for each positive winding number $n\geq 1$ as
$\alpha(\theta)$ $=$ $\alpha\left(\theta^{0}_{n\pm}\right)+\left.\frac{\mathrm{d}\alpha}{\mathrm{d}\theta}\right|_{\theta=\theta^{0}_{n\pm}}\left(\theta-\theta^{0}_{n\pm}\right)$ (72)
$+O\left(\left(\theta-\theta^{0}_{n\pm}\right)^{2}\right).$
From Eqs. (64), (68), (70), and (72), the effective deflection angle $\bar{\alpha}\left(\theta_{n\pm}\right)$, where $\theta=\theta_{n\pm}$ is the image angle for each positive winding number $n\geq 1$, is given by
$\bar{\alpha}\left(\theta_{n\pm}\right)=\mp\frac{\bar{c}_{\pm}\left(\theta_{n\pm}-\theta^{0}_{n\pm}\right)}{6\theta_{\infty}\left|\frac{\theta^{0}_{n\pm}}{\theta_{\infty}}-1\right|^{7/6}}.$ (73)
From Eqs. (65), (71), and (73), the image angles $\theta_{n\pm}=\theta_{n\pm}(\phi)$ of the rays reflected inside and outside of the marginally unstable photon sphere for each positive winding number $n\geq 1$ are obtained as
$\theta_{n\pm}(\phi)\sim\theta^{0}_{n\pm}\mp\frac{d_{\mathrm{OS}}}{d_{\mathrm{LS}}}\frac{6\bar{c}_{\pm}^{6}}{\left(2\pi n-\bar{d}_{\pm}\right)^{7}}\theta_{\infty}(\theta^{0}_{n\pm}-\phi).$ (74)
If the observer O, the lens L, and the source S are aligned in a straight line in this order, the observer sees ring-shaped images called Einstein rings. By setting $\phi=0$, we obtain the image angle of the Einstein ring for each positive winding number $n\geq 1$ as
$\theta_{\mathrm{E}n\pm}\equiv\theta_{n\pm}(0)\sim\theta^{0}_{n\pm}\left[1\mp\frac{d_{\mathrm{OS}}}{d_{\mathrm{LS}}}\frac{6\bar{c}_{\pm}^{6}}{\left(2\pi n-\bar{d}_{\pm}\right)^{7}}\theta_{\infty}\right].$ (75)
The difference of the image angles between the image with $n=1$ and the marginally unstable photon sphere with $n=\infty$ are obtained as
$\bar{s}_{\pm}\equiv\theta_{1\pm}-\theta_{\infty}\sim\theta^{0}_{1\pm}-\theta^{0}_{\infty}=\pm\left(\frac{\bar{c}_{\pm}}{2\pi-\bar{d}_{\pm}}\right)^{6}\theta_{\infty}.$ (76)
The magnifications of the images of the ray reflected by just inside and outside of the marginally unstable photon sphere for each positive winding number $n\geq 1$ are given by
$\mu_{n\pm}(\phi)$ $\equiv$ $\frac{\theta_{n\pm}}{\phi}\frac{\mathrm{d}\theta_{n\pm}}{\mathrm{d}\phi}$ (77)
$\sim$ $\frac{d_{\mathrm{OS}}}{d_{\mathrm{LS}}}\frac{6\bar{c}_{\pm}^{6}\left[\bar{c}_{\pm}^{6}\pm(2\pi n-\bar{d}_{\pm})^{6}\right]}{(2\pi n-\bar{d}_{\pm})^{13}}\frac{\theta_{\infty}^{2}}{\phi}.$
Every gravitational lensed image with the winding number $n\geq 1$ and with a positive impact parameter makes a pair of the image with the same winding number $n\geq 1$ and with a negative impact parameter. The pair of images with every winding number $n\geq 1$ has the almost same absolute values of the image angles and the magnifications but a different signs under the assumptions of the small angles. Thus, the diameter of the paired images is given by $2\theta_{n\pm}$ with the winding number $n\geq 1$ for the positive impact parameter and the total magnification $\mu_{n\mathrm{tot}\pm}$, which is defined by sum of the absolute value of the magnification of the pair of the images with the winding number $n\geq 1$, is given by
$\mu_{n\mathrm{tot}\pm}\sim 2\left|\mu_{n\pm}\right|.$ (78)
The ratio of the magnification of the image with $n=1$ to the sum of the other images with $n\geq 2$ is given by
$\bar{r}_{\pm}\equiv\frac{\mu_{1\pm}}{\sum^{\infty}_{n=2}\mu_{n\pm}}\sim\frac{K_{1\pm}}{\sum^{\infty}_{n=2}K_{n\pm}},$ (79)
where $K_{n\pm}$ for each positive winding number $n\geq 1$ is defined by
$K_{n\pm}\equiv\frac{\bar{c}_{\pm}^{6}\pm(2\pi n-\bar{d}_{\pm})^{6}}{(2\pi n-\bar{d}_{\pm})^{13}}.$ (80)
In Table I, the observable with the distances $d_{\mathrm{OS}}=16$kpc and $d_{\mathrm{OL}}=d_{\mathrm{LS}}=8$kpc, the mass $M=4\times 10^{6}M_{\odot}$ in the Reissner-Nordström spacetime and the Hayward spacetime is shown for the winding numbers $n=1$, 22, and $3$.
Table 1: The coefficients $\bar{c}_{\pm}$ and the terms $\bar{d}_{\pm}$ in the deflection angles of the rays deflected inside and outside of the marginally unstable photon spheres in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ in the Reissner-Nordström (RN) spacetime obtained in our numerical method, by Sasaki Sasaki:2025web , and a semi-analytical method by Tsukamoto Tsukamoto:2020iez and the ones in the Hayward spacetime in our method, by Chiba and Kimura Chiba:2017nml , and in the semi-analytical method by Tsukamoto Tsukamoto:2020iez , and observable calculated by them are shown in the case with distances $d_{\mathrm{OS}}=16$kpc and $d_{\mathrm{OL}}=d_{\mathrm{LS}}=8$kpc, and the mass $M=4\times 10^{6}M_{\odot}$. The reduced critical impact parameter $b_{\mathrm{m}}/M$, the diameter of the marginally unstable photon sphere $2\theta_{\infty}$, the diameters of the Einstein rings $2\theta_{\mathrm{E}1\pm}$, $2\theta_{\mathrm{E}2\pm}$, and $2\theta_{\mathrm{E}3\pm}$ with the winding number $n=1$, 22, and $3$, respectively, the difference of the radii $\bar{s}_{\pm}=\theta_{\mathrm{E}1\pm}-\theta_{\infty}$, the total magnifications of the pair images $\mu_{1\mathrm{tot}\pm}(\phi)$, $\mu_{2\mathrm{tot}\pm}(\phi)$, and $\mu_{3\mathrm{tot}\pm}(\phi)$ with the source angle $\phi=1$ arcsecond, and with $n=1$, 22, and $3$, respectively, and the ratio of the magnifications $\bar{r}_{\pm}$ are shown.
RN in our method RN in Sasaki:2025web RN in Tsukamoto:2020iez Hayward in our method Hayward in Chiba:2017nml Hayward in Tsukamoto:2020iez
$b_{\mathrm{m}}/M$ $3\sqrt{6}/2$ $3\sqrt{6}/2$ $3\sqrt{6}/2$ $25\sqrt{5}/12$ $25\sqrt{5}/12$ $25\sqrt{5}/12$
[…]
$\bar{c}_{+}$ 6.67748 6.67748 5.49892 6.01322 6.01316 4.95196
$\bar{d}_{+}$ -5.59011 -5.59108 -5.59108 -5.60584 - -5.62607
[…]
$\mu_{1\mathrm{tot}+}(\phi)\times 10^{19}$ 1063.82 1063.19 324.595 890.522 - 271.347
$\mu_{2\mathrm{tot}+}(\phi)\times 10^{19}$ 52.8675 52.8477 16.4542 44.996 - 13.9129
$\mu_{3\mathrm{tot}+}(\phi)\times 10^{19}$ 6.58948 6.58765 2.05397 5.62252 - 1.7433
$\bar{r}_{+}$ 17.3389 17.3350 16.9947 17.0469 - 16.7917
$\bar{c}_{-}$ 11.5658 11.5657 - 10.4155 - -
$\bar{d}_{-}$ -5.64071 -5.59108 - -5.69675 - -
[…]
$\mu_{1\mathrm{tot}-}(\phi)\times 10^{19}$ 4518.59 4066.21 - 12739.8 - -
$\mu_{2\mathrm{tot}-}(\phi)\times 10^{19}$ 1304.72 1328.28 - 1131.53 - -
$\mu_{3\mathrm{tot}-}(\phi)\times 10^{19}$ 173.347 175.793 - 147.045 - -
$\bar{r}_{-}$ 2.95597 2.61464 - 9.64064 - -
V Discussion and Conclusion
In this paper, we have extent Eiroa, Romero, and Torres’s method Eiroa:2002mk to gravitational lensing of rays just inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and we have applied it to a Reissner-Nordström spacetime and a Hayward spacetime with the marginally unstable photon sphere.
We have confirmed that the deflection angles in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ by the method converge correctly to the deflection angle without approximations, while there were mismatches in semi-analytic calculations by the author previously Tsukamoto:2020iez , which was pointed out by Sasaki Sasaki:2025web , for the coefficient $\bar{c}_{+}$ of the power-divergent term of the deflection angles of the rays deflected just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime.
In Ref. Sasaki:2025web , Sasaki found that the term $\bar{d}_{-}$ has the same as the term $\bar{d}_{+}$ analytically in the Reissner-Nordström spacetime with the marginally unstable photon sphere. We obtained $\bar{d}_{+}\sim-5.59011$ and $\bar{d}_{-}\sim-5.64071$ as shown Eq. (46) and Eq. (48), respectively, and its deference is few percentages by our method. We note that we have approximately read $r_{0}\rightarrow r_{\mathrm{m}}+0$ in Eqs. (28) and (29) as $r_{0}=r_{\mathrm{m}}+10^{-7}M$ for numerical calculations to get the values of $\bar{c}_{+}$ and $\bar{d}_{+}$ and we have used $r_{0}=r_{\mathrm{m}}-10^{-5}M$ for numerical calculations of $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs. (28) and (29) to get the values of $\bar{c}_{-}$ and $\bar{d}_{-}$. One would reduce numerical errors by calculating the integrals for the deflection angles more precisely.
If the relation $\bar{d}_{-}=\bar{d}_{+}$ is also correct in other spacetimes, the relation would be used to check our numerical method. In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{d}_{+}\sim-5.60584$ and $\bar{d}_{-}\sim-5.69675$ in our method and they imply that the relation $\bar{d}_{-}=\bar{d}_{+}$ is almost held. We have approximately read $r_{0}\rightarrow r_{\mathrm{m}}+0$ in Eqs. (28) and (29) as $r_{0}=r_{\mathrm{m}}+5\times 10^{-5}M$ for numerical calculations to get the values of $\bar{c}_{+}$ and $\bar{d}_{+}$ and we have used $r_{0}=r_{\mathrm{m}}-5\times 10^{-5}M$ for numerical calculations of $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs. (28) and (29) to get the values of $\bar{c}_{-}$ and $\bar{d}_{-}$.
In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{c}_{+}\sim 6.01322$ in our numerical method and its difference from the value of $\bar{c}_{+}\sim 6.01316$ in Chiba and Kimura Chiba:2017nml is tiny but we have found that the value $\bar{c}_{+}\sim 4.95196$ obtained by the semi-analytic calculation in Tsukamoto:2020iez should be modified.
Thus, we conclude that our extent Eiroa, Romero, and Torres’s method to gravitational lensing of the rays just inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ gives the correct values of $\bar{c}_{\pm}$ and $\bar{d}_{\pm}$ and that the semi-analytic method considered by the author previously in Ref. Tsukamoto:2020iez gives the invalid values of $\bar{c}_{+}$ while it gives the correct values of $\bar{d}_{+}$.
We comment on constraints on metrics from the shadow observations of supermassive black hole candidates at the centers of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp . The Event Horizon Telescope collaborations EventHorizonTelescope:2021dqv ; EventHorizonTelescope:2022xqj constrained the parameter of the metrics of spacetimes from the observations and the sizes of the photon spheres under assumptions for simplify. Thus, the constraints. which depend on the assumptions. are not very reliable Gralla:2020pra ; Tsukamoto:2024gkz and the constraints on the metrics should be confirmed by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr .

### v2

@@
 It is believed that rays bent inside and outside photon spheres could affect partially the black hole shadow images by the Event Horizon Telescope and the rays near photon spheres would be detected by near-future space observations.
-The features of the rays near the photon spheres in not only black hole spacetimes but also exotic spacetimes would be important since one will need them to exclude black-hole mimickers.
-The deflection angles of the rays deflected by photon spheres diverge logarithmically and we can treat them by a strong-deflection-limit analysis.
+The investigation of the rays near the photon spheres in not only black hole spacetimes but also exotic spacetimes would be important since one will need them to exclude black-hole mimickers.
+The deflection angles of the rays deflected by the photon spheres diverge logarithmically and we can treat them by a strong-deflection-limit analysis.
 The error of the strong-deflection-limit analysis becomes large if antiphoton spheres exist in the spacetimes and the analysis breaks down when the photon spheres and the antiphoton spheres degenerate to form a marginally unstable photon sphere.
@@
 The gravitational lenses have been mainly studied in a weak gravitational field for a hundred years since we do not have enough technology to detect astrophysical phenomena in strong gravitational fields until the last decade.
-Rays deflected by the photon spheres of the black holes and their gravitational lensing effects in strong gravitational fields have been intermittently investigated Hagihara_1931 ; Darwin_1959 ; Atkinson_1965 ; Luminet_1979 ; Ohanian_1987 ; Nemiroff_1993 ; Frittelli_Kling_Newman_2000 ; Virbhadra_Ellis_2000 ; Bozza:2001xd ; Bozza:2002zj ; Perlick:2003vg ; Nandi:2006ds ; Virbhadra:2008ws ; Hioki:2009na ; Bozza_2010 ; Tsupko:2017rdo ; Nakao:2018knn ; Gralla:2019xty ; Okabayashi:2020apz ; Hsieh:2021scb ; Aratore:2021usi ; Tsupko:2022kwi ; Aratore:2024bro ; Tsupko:2025hhf .
+Rays deflected by the photon spheres of the black holes and their gravitational lensing effects in strong gravitational fields have been intermittently investigated Hagihara_1931 ; Darwin_1959 ; Atkinson_1965 ; Luminet_1979 ; Ohanian_1987 ; Nemiroff_1993 ; Frittelli_Kling_Newman_2000 ; Virbhadra_Ellis_2000 ; Bozza:2001xd ; Bozza:2002zj ; Perlick:2003vg ; Nandi:2006ds ; Virbhadra:2008ws ; Hioki:2009na ; Bozza_2010 ; Tsupko:2017rdo ; Nakao:2018knn ; Gralla:2019xty ; Okabayashi:2020apz ; Hsieh:2021scb ; Aratore:2021usi ; Tsupko:2022kwi ; Aratore:2024bro ; Feleppa:2024vdk ; Feleppa:2025ejh ; Tsupko:2025hhf ; Saketh:2025cwf .
 The effects of the photon spheres on the apparent shape of stars collapsing into the black holes also have been investigated Ames_1968 ; Synge:1966okc ; Yoshino:2019qsh ; Koga:2025dii .
@@
 It is important to investigate the features of the rays near the photon spheres in not only black hole spacetimes but also exotic compact objects such as wormholes Ellis:1973yv ; Chetouani:1984qdm ; Perlick:2003vg ; Muller:2004dq ; Nandi:2006ds ; Muller:2008zza ; Tsukamoto:2012xs ; Ohgami:2015nra ; Tsukamoto:2016qro ; Tsukamoto:2016zdu ; Shaikh:2018kfv ; Shaikh:2018oul ; Bronnikov:2018nub ; Shaikh:2019jfr ; Bronnikov:2021liv ; Maeda:2021wnl ; Olmo:2023lil ; Bugaev:2023mlc ; Zhang:2024sgs ; Xavier:2024iwr ; Tsukamoto:2024pid ; Bugaev:2024dte ; Solodukhin:2025opw and naked singularity Virbhadra:1998dy ; Virbhadra:2002ju ; Hioki:2009na ; Shaikh:2018lcc ; Shaikh:2019itn since one needs to exclude exotic compact objects or black-hole mimickers to reinforce claims that the observed objects must be supermassive black holes.
-Gravitational lensing of rays reflected by just outside of a photon sphere in a Schwarzschild spacetime in a strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$, where $b$ and $b_{\mathrm{m}}$ an impact parameter and a critical impact parameter of the ray, by Bozza et al.
+Gravitational lensing of rays reflected by just outside of a photon sphere in a Schwarzschild spacetime in a strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$, where $b$ and $b_{\mathrm{m}}$ are the impact parameter and the critical impact parameter of the ray, by Bozza et al.
 Bozza:2001xd .
@@
 The strong-deflection-limits analysis has been intensively studied and it has been extended over the last two decades  Bozza:2002af ; Eiroa:2002mk ; Petters:2002fa ; Eiroa:2003jf ; Bozza:2004kq ; Bozza:2005tg ; Bozza:2006sn ; Bozza:2006nm ; Iyer:2006cn ; Bozza:2007gt ; Tsukamoto:2016qro ; Ishihara:2016sfv ; Tsukamoto:2016oca ; Tsukamoto:2016zdu ; Tsukamoto:2016jzh ; Shaikh:2018oul ; Shaikh:2019itn ; Takizawa:2021gdp ; Zhang:2024sgs ; Igata:2025taz .
-Tsukamoto Tsukamoto:2016oca ; Tsukamoto:2016jzh investigated an alternative of Bozza’s method Bozza:2002zj , obtained $\bar{a}_{+}$ and $\bar{b}_{+}$ in analytic forms in the Reissner-Nordström spacetime and confirmed that they are consistent with $\bar{a}_{+}$ and $\bar{b}_{+}$ obtained by Eiroa, Romero, and Torres Eiroa:2002mk and by Bozza Bozza:2002zj .
+Tsukamoto Tsukamoto:2016oca ; Tsukamoto:2016jzh investigated an alternative of Bozza’s method Bozza:2002zj , obtained $\bar{a}_{+}$ and $\bar{b}_{+}$ in analytic forms in the Reissner-Nordström black hole spacetime and confirmed that they are consistent with $\bar{a}_{+}$ and $\bar{b}_{+}$ obtained by Eiroa, Romero, and Torres Eiroa:2002mk and by Bozza Bozza:2002zj for the Reissner-Nordström black hole spacetime.
 If we consider gravitational lensing effects of rays passing inside and outside of the photon spheres in spacetimes without event horizons Shaikh:2019itn ; Shaikh:2019jfr , the deflection angles of the rays in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ can be expressed by
@@
 In this paper, we extend Eiroa, Romero, and Torres’s method Eiroa:2002mk to gravitational lensing of rays just inside and outside of the marginally unstable photon sphere in the general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and we apply it to the Reissner-Nordström spacetime and a Hayward spacetime Hayward:2005gi with the marginally unstable photon sphere.
-We also confirm that mismatches in the coefficients $\bar{c}_{+}$ which were calculated by the author in the semi-analytic method previously Tsukamoto:2020iez , are modified in the new approach.
+We also confirm that mismatches in the coefficients $\bar{c}_{+}$ which were calculated by Tsukamoto in the semi-analytic method previously Tsukamoto:2020iez , are modified in the new approach.
 This paper is organized as follows.
@@
 We conclude our results in Sec.
-V.
+V and discuss in Sec.
+VI.
 We use a unit that a light speed and Newton’s constant are unity in this paper.
@@
 $D(r)\equiv\frac{C^{\prime}(r)}{C(r)}-\frac{A^{\prime}(r)}{A(r)},$ (8)
-where denotes a differentiation with respect to $r$ and functions with the subtract m denotes the functions at $r=r_{\mathrm{m}}$. The circular light orbits are unstable (stable) if a condition
+where denotes a differentiation with respect to $r$ and functions with the subscript m denotes the functions at $r=r_{\mathrm{m}}$. The circular light orbits are unstable (stable) if a condition
 $D_{\mathrm{m}}^{\prime}=\frac{C_{\mathrm{m}}^{\prime\prime}}{C_{\mathrm{m}}}-\frac{A^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}>0\quad(<0)$ (9)
@@
 The spacetime has time-translational and axial Killing vectors $t^{\mu}\partial_{\mu}=\partial_{t}$ and $\varphi^{\mu}\partial_{\mu}=\partial_{\varphi}$ due to stationarity and axial symmetry of the spacetime, respectively, and we can assume $\vartheta=\pi/2$ without loss of generality because of the spherical symmetry of the spacetime.
-From the wave number of a ray $k^{\mu}\equiv\dot{x}^{\mu}$, where $x^{\mu}$ is the coordinates and the dot denotes a differentiation with respect to an affine parameter along the trajectory of the ray, the trajectory of the ray is obtained as
+From
+$k^{\mu}k_{\mu}=0,$ (12)
+where $k^{\mu}$ is the wave number of a ray defined as $k^{\mu}\equiv\dot{x}^{\mu}$, where $x^{\mu}$ is the coordinates and the dot denotes a differentiation with respect to an affine parameter along the trajectory of the ray, the trajectory of the ray is obtained as
 $A(r)\dot{t}^{2}+B(r)\dot{r}^{2}+C(r)\dot{\varphi}^{2}=0.$ (13)
 We consider that the ray, which comes from a spatial infinity, is reflected by a lensing object at the closest distance $r=r_{0}$ and is away to the spatial infinity.
-In this paper, we do not consider a case that the rays coming from the spatial infinity, passes through a wormhole throat, and leave to another spatial infinity if we consider wormhole spacetimes.
+In this paper, we do not consider a case that the rays coming from the spatial infinity, passes through a wormhole throat, and leave to another spatial infinity if we consider a wormhole spacetime.
 At the closest distance $r=r_{0}$, due to
 $\dot{r}_{0}\equiv\left.\dot{r}\right|_{r=r_{0}}=0,$ (14)
-where functions with the subtract 0 denote the functions at the closest distance $r=r_{0}$, we obtain, from Eq. (12),
+where functions with the subscript 0 denote the functions at the closest distance $r=r_{0}$, we obtain, from Eq. (13),
 $A_{0}\dot{t}_{0}^{2}=C_{0}\dot{\varphi}^{2}_{0}.$ (15)
 From Eq.
-(14), the impact parameter $b$ of the ray defined as $b\equiv L/E$, where $E\equiv-g_{\mu\nu}t^{\mu}k^{\nu}$ and $L\equiv g_{\mu\nu}\varphi^{\mu}k^{\nu}$ are the conserved energy and the angular momentum of the ray, respectively, is expressed as
+(15), the impact parameter $b$ of the ray defined as $b\equiv L/E$, where $E\equiv-g_{\mu\nu}t^{\mu}k^{\nu}$ and $L\equiv g_{\mu\nu}\varphi^{\mu}k^{\nu}$ are the conserved energy and angular momentum of the ray, respectively, is expressed as
 $b=b(r_{0})=\frac{L}{E}=\frac{C_{0}\dot{\varphi}_{0}}{A_{0}\dot{t}_{0}}=\sqrt{\frac{C_{0}}{A_{0}}},$ (16)
@@
 $D^{\prime\prime}_{\mathrm{m}}=\frac{C^{\prime\prime\prime}_{\mathrm{m}}}{C_{\mathrm{m}}}-\frac{A^{\prime\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}}>0$ (23)
-is satisfied so that the spacetime has the marginally unstable photon sphere. Under the assumptions, we obtain
+are satisfied so that the spacetime has the marginally unstable photon sphere. Under the assumptions, we obtain
 $V(r_{\mathrm{m}},b_{\mathrm{m}})=V^{\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=V^{\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=0,$ (24)
-$V^{\prime\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=-\frac{b_{\mathrm{m}}D^{\prime\prime}_{\mathrm{m}}}{B_{\mathrm{m}}C_{\mathrm{m}}}=-\frac{D^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}B_{\mathrm{m}}}<0,$ (24)
-and the deflection angle of the ray bent outside and inside of the marginally unstable photon sphere in the strong deflection limits $r_{0}\rightarrow r_{\mathrm{m}}\pm 0$ or $b\rightarrow b_{\mathrm{m}}\pm 0$ will be expressed by,
+$V^{\prime\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=-\frac{b_{\mathrm{m}}^{2}D^{\prime\prime}_{\mathrm{m}}}{B_{\mathrm{m}}C_{\mathrm{m}}}=-\frac{D^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}B_{\mathrm{m}}}<0,$ (25)
+and the deflection angles of the rays bent outside and inside of the marginally unstable photon sphere in the strong deflection limits $r_{0}\rightarrow r_{\mathrm{m}}\pm 0$ or $b\rightarrow b_{\mathrm{m}}\pm 0$ will be expressed by,
 $\alpha=\frac{c_{\pm}}{\left|r_{0}-r_{\mathrm{m}}\right|^{1/2}}+\bar{d}_{\pm},$ (26)
-where $c_{\pm}$ is the coefficient and $\bar{d}_{\pm}$ is the term. From the relation
+where $c_{\pm}$ are coefficients and $\bar{d}_{\pm}$ are finite terms. From the relation
 $\lim_{r_{0}\rightarrow r_{\mathrm{m}}\pm 0}\left[\alpha-\frac{c_{\pm}}{\left|r_{0}-r_{\mathrm{m}}\right|^{1/2}}-\bar{d}_{\pm}\right]=0,$ (27)
@@
 Aratore:2024bro ; Tsupko:2025hhf by using a strong-deflection-limit analysis with arbitrary source distances Bozza:2007gt .
-In the Reissner-Nordström naked singularity spacetime, the size of the photon sphere or the shadows Zakharov:2014lqa and gravitational lensing of rays deflected inside and outside of a photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}\pm 0$ were investigated in Refs.
+In the Reissner-Nordström naked singularity spacetime, the size of the photon sphere or the shadows Zakharov:2014lqa and gravitational lensing of rays deflected inside and outside of a photon sphere in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ were investigated in Refs.
 Shaikh:2019itn ; Tsukamoto:2021fsz ; Tsukamoto:2021lpm .
@@
 The broken (green), solid (red), and dotted (magenta) curves denote $V(r/M,b)$ for $b=0.99b_{\mathrm{m}}$, $b_{\mathrm{m}}$, and $1.01b_{\mathrm{m}}$, respectively.
-From Eqs. (28), (29), and (32), we obtain
+From Eqs. (29), (30), and (33), we obtain numerically
 $\bar{c}_{+}\sim 6.67748$ (46)
@@
 $\bar{d}_{+}\sim-5.59011$ (47)
-numerically for the rays just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime.
-From Eqs. (28) and (29), for the rays just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime, we get
+for the rays just outside of the marginally unstable photon sphere, and
 $\bar{c}_{-}\sim 11.5658$ (48)
@@
 $\bar{d}_{-}\sim-5.64071$ (49)
-in numerical.
-The deflection angle $\alpha$ is shown in Fig. 2 and the percent error of the deflection angle of Eq. (3.12) defined by
+for the ones inside of it in the Reissner-Nordström spacetime.
+The deflection angles $\alpha$ are shown in Fig. 2 and the percent error of the deflection angle of Eq. (3.12) defined by
 $\frac{\alpha\>\mathrm{of\>Eq.}(3.12)-\alpha\>\mathrm{of\>Eq.}(2.17)}{\alpha\>\mathrm{of\>Eq.}(2.17)}\times 100$ (50)
-is shown in Fig. 3.
+is shown in Fig.
+3.
+From Figs.
+2 and 3, we confirm that our results are the almost same as ones by Sasaki Sasaki:2025web and that they converge to the deflection angle (2.17) in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$.
+At a glance at Fig.
+2, it might seem as the deflection angle by Tsukamoto Tsukamoto:2020iez converge to the deflection angle (2.17) in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ but Fig.
+3 shows that it does not converge as pointed out by Sasaki Sasaki:2025web .
 Figure 2: The deflection angles $\alpha$ in the Reissner-Nordström spacetime by using Eq.
@@
 where
-$c_{+}=2^{11/6}3^{2/3}5^{5/4}\int^{\infty}_{0}\frac{dy}{\sqrt{432y+900y^{2}+625y^{3}}}\sim 7.771.$ (51)
+$c_{+}=2^{11/6}3^{2/3}5^{5/4}\int^{\infty}_{0}\frac{\mathrm{d}y}{\sqrt{432y+900y^{2}+625y^{3}}}\sim 7.771.$ (52)
 but they do not consider the constants $\bar{d}_{+}$ and $\bar{d}_{-}$ and the coefficient $\bar{c}_{-}$.
-Tsukamoto Tsukamoto:2020iez studied the deflection angle of the rays reflected outside of a marginally unstable photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ and got the coefficient as, by using the semi-analytic method,
+Tsukamoto Tsukamoto:2020iez studied the deflection angle of the rays reflected outside of the marginally unstable photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ and got the coefficient as, by using the semi-analytic method,
 $\bar{c}_{+}\sim\frac{2^{13/6}3^{1/3}}{5^{1/6}}\sim 4.95196$ (53)
 and the constant
-$\bar{d}_{+}=-4\sqrt{\frac{6}{5}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi\sim-5.62607.$ (53)
-The percent errors of the deflection angle of Eq. (3.12) with the coefficients $\bar{c}_{+}$ and the constants $\bar{d}_{+}$ are plotted in Fig. 6.
+$\bar{d}_{+}=-4\sqrt{\frac{6}{5}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi\sim-5.62607,$ (54)
+where an integral
+$I_{\mathrm{R}}(r_{\mathrm{m}})=2\int^{1}_{0}\left[\sqrt{\frac{6-3z+3z^{2}-z^{3}}{z^{3}(5-5z+z^{2})}}-\sqrt{\frac{6}{5z^{3}}}\right]\mathrm{d}z$ (55)
+was calculated in numerical.
 The Hayward spacetime Hayward:2005gi has the marginally unstable photon sphere at $r=r_{\mathrm{m}}=25M/12$ made by the rays with the critical impact parameter
@@
 Figure 4: The effective potentials $V(r/M,b)$ for $b=0.99b_{\mathrm{m}}$, $b_{\mathrm{m}}$, and $1.01b_{\mathrm{m}}$ in the Hayward spacetime with the marginally unstable photon sphere are shown as the broken (green), solid (red), and dotted (magenta) curves, respectively.
-From Eqs. (28), (29), and (32), we obtain
-$\bar{c}_{+}\sim 6.01322$ (59)
+From Eqs. (29), (30), and (33), we obtain in numerical
+$\bar{c}_{+}\sim 6.01324$ (61)
 and
-$\bar{d}_{+}\sim-5.60584$ (60)
-in numerical for the Hayward spacetime with the marginally unstable photon sphere. From Eqs. (28) and (29), we get the numerical values of
-$\bar{c}_{-}\sim 10.4155$ (61)
+$\bar{d}_{+}\sim-5.60958$ (62)
+for the rays outside of the marginally unstable photon sphere
+$\bar{c}_{-}\sim 10.4154$ (63)
 and
-$\bar{d}_{-}\sim-5.69675.$ (62)
-The deflection angle $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere is shown in Fig.
-5 and the percent error of the deflection angle of Eq.
-(3.12) with the coefficients $\bar{c}_{+}$ and the constants $\bar{d}_{+}$ is shown in Fig.
+$\bar{d}_{-}\sim-5.67374.$ (64)
+for the ones inside of the marginally unstable photon sphere in the Hayward spacetime.
+The deflection angles $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere are shown in Fig.
+5 and the percent errors of the deflection angles of Eq.
+(3.12) with the coefficients $\bar{c}_{\pm}$ and the constants $\bar{d}_{\pm}$ are shown in Fig.
 6.
+From Fig.
+5, one may consider that the deflection angle by Tsukamoto Tsukamoto:2020iez converges to the deflection angle (2.17) in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ but Fig.
+6 shows that it does not converge.
+From Fig.
+6, we find that the deflection angle which we obtained in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and one by Chiba and Kimura Chiba:2017nml in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ converges to the deflection angle (2.17) and we notice that the errors of our calculations are smaller than one by Chiba and Kimura Chiba:2017nml due to their lack of the constant $\bar{d}_{+}$.
+Therefore, we recognize that the coefficient $\bar{c}_{+}$ (51) by Chiba and Kimura Chiba:2017nml and the constant $\bar{d}_{+}$ (55) by Tsukamoto Tsukamoto:2020iez are correct while the coefficient $\bar{c}_{+}$ (53) by using the semi-analytic method in Ref.Tsukamoto:2020iez should be modified.
 Figure 5: The deflection angles $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere are shown.
@@
 (2.17) in the Hayward spacetime are plotted in top and bottom panels for the rays producing images outside and inside of the marginally unstable photon sphere, respectively.
-In the top pannel, the dashed (green), dotted (black), and solid (red) curves denote the percent errors of deflection angles $\alpha$ of Eq.
+In the top pannel, the dashed (green), dotted (black), and solid (red) curves denote the percent errors of the deflection angles $\alpha$ of Eq.
 (3.12) with $\bar{c}_{+}$ of Eq.
@@
 (64) in our numerical method.
-We have confirmed the coefficient $\bar{c}_{+}$ (50) by Chiba and Kimura Chiba:2017nml and the constant $\bar{d}_{+}$ (53) by Tsukamoto Tsukamoto:2020iez are correct while the coefficient $\bar{c}_{+}$ (52) by using the semi-analytic method in Ref.Tsukamoto:2020iez should be modified.
 IV Observable of gravitational lensing in the strong deflection limits
@@
 The pair of images with every winding number $n\geq 1$ has the almost same absolute values of the image angles and the magnifications but a different signs under the assumptions of the small angles.
-Thus, the diameter of the paired images is given by $2\theta_{n\pm}$ with the winding number $n\geq 1$ for the positive impact parameter and the total magnification $\mu_{n\mathrm{tot}\pm}$, which is defined by sum of the absolute value of the magnification of the pair of the images with the winding number $n\geq 1$, is given by
+Thus, the diameter of the paired images is given by $2\theta_{n\pm}$ with the winding number $n\geq 1$ for the positive impact parameter and the total magnification $\mu_{n\mathrm{tot}\pm}$, which is defined by the sum of the absolute values of the magnification of the pair of the images with the winding number $n\geq 1$, is given by
 $\mu_{n\mathrm{tot}\pm}\sim 2\left|\mu_{n\pm}\right|.$ (80)
@@
 $K_{n\pm}\equiv\frac{\bar{c}_{\pm}^{6}\pm(2\pi n-\bar{d}_{\pm})^{6}}{(2\pi n-\bar{d}_{\pm})^{13}}.$ (82)
-In Table I, the observable with the distances $d_{\mathrm{OS}}=16$kpc and $d_{\mathrm{OL}}=d_{\mathrm{LS}}=8$kpc, the mass $M=4\times 10^{6}M_{\odot}$ in the Reissner-Nordström spacetime and the Hayward spacetime is shown for the winding numbers $n=1$, 22, and $3$.
-Table 1: The coefficients $\bar{c}_{\pm}$ and the terms $\bar{d}_{\pm}$ in the deflection angles of the rays deflected inside and outside of the marginally unstable photon spheres in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ in the Reissner-Nordström (RN) spacetime obtained in our numerical method, by Sasaki Sasaki:2025web , and a semi-analytical method by Tsukamoto Tsukamoto:2020iez and the ones in the Hayward spacetime in our method, by Chiba and Kimura Chiba:2017nml , and in the semi-analytical method by Tsukamoto Tsukamoto:2020iez , and observable calculated by them are shown in the case with distances $d_{\mathrm{OS}}=16$kpc and $d_{\mathrm{OL}}=d_{\mathrm{LS}}=8$kpc, and the mass $M=4\times 10^{6}M_{\odot}$.
-The reduced critical impact parameter $b_{\mathrm{m}}/M$, the diameter of the marginally unstable photon sphere $2\theta_{\infty}$, the diameters of the Einstein rings $2\theta_{\mathrm{E}1\pm}$, $2\theta_{\mathrm{E}2\pm}$, and $2\theta_{\mathrm{E}3\pm}$ with the winding number $n=1$, 22, and $3$, respectively, the difference of the radii $\bar{s}_{\pm}=\theta_{\mathrm{E}1\pm}-\theta_{\infty}$, the total magnifications of the pair images $\mu_{1\mathrm{tot}\pm}(\phi)$, $\mu_{2\mathrm{tot}\pm}(\phi)$, and $\mu_{3\mathrm{tot}\pm}(\phi)$ with the source angle $\phi=1$ arcsecond, and with $n=1$, 22, and $3$, respectively, and the ratio of the magnifications $\bar{r}_{\pm}$ are shown.
-RN in our method RN in Sasaki:2025web RN in Tsukamoto:2020iez Hayward in our method Hayward in Chiba:2017nml Hayward in Tsukamoto:2020iez
-$b_{\mathrm{m}}/M$ $3\sqrt{6}/2$ $3\sqrt{6}/2$ $3\sqrt{6}/2$ $25\sqrt{5}/12$ $25\sqrt{5}/12$ $25\sqrt{5}/12$
-[…]
-$\bar{c}_{+}$ 6.67748 6.67748 5.49892 6.01322 6.01316 4.95196
-$\bar{d}_{+}$ -5.59011 -5.59108 -5.59108 -5.60584 - -5.62607
-[…]
-[…]
-[…]
-[…]
-$\mu_{1\mathrm{tot}+}(\phi)\times 10^{19}$ 1063.82 1063.19 324.595 890.522 - 271.347
-$\mu_{2\mathrm{tot}+}(\phi)\times 10^{19}$ 52.8675 52.8477 16.4542 44.996 - 13.9129
-$\mu_{3\mathrm{tot}+}(\phi)\times 10^{19}$ 6.58948 6.58765 2.05397 5.62252 - 1.7433
-$\bar{r}_{+}$ 17.3389 17.3350 16.9947 17.0469 - 16.7917
-$\bar{c}_{-}$ 11.5658 11.5657 - 10.4155 - -
-$\bar{d}_{-}$ -5.64071 -5.59108 - -5.69675 - -
-[…]
-[…]
-[…]
-[…]
-$\mu_{1\mathrm{tot}-}(\phi)\times 10^{19}$ 4518.59 4066.21 - 12739.8 - -
-$\mu_{2\mathrm{tot}-}(\phi)\times 10^{19}$ 1304.72 1328.28 - 1131.53 - -
-$\mu_{3\mathrm{tot}-}(\phi)\times 10^{19}$ 173.347 175.793 - 147.045 - -
-$\bar{r}_{-}$ 2.95597 2.61464 - 9.64064 - -
-V Discussion and Conclusion
+V Conclusion
 In this paper, we have extent Eiroa, Romero, and Torres’s method Eiroa:2002mk to gravitational lensing of rays just inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and we have applied it to a Reissner-Nordström spacetime and a Hayward spacetime with the marginally unstable photon sphere.
@@
 If the relation $\bar{d}_{-}=\bar{d}_{+}$ is also correct in other spacetimes, the relation would be used to check our numerical method.
-In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{d}_{+}\sim-5.60584$ and $\bar{d}_{-}\sim-5.69675$ in our method and they imply that the relation $\bar{d}_{-}=\bar{d}_{+}$ is almost held.
+In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{d}_{+}\sim-5.60958$ and $\bar{d}_{-}\sim-5.67374$ in our method and they imply that the relation $\bar{d}_{-}=\bar{d}_{+}$ is almost held.
 We have approximately read $r_{0}\rightarrow r_{\mathrm{m}}+0$ in Eqs.
-(28) and (29) as $r_{0}=r_{\mathrm{m}}+5\times 10^{-5}M$ for numerical calculations to get the values of $\bar{c}_{+}$ and $\bar{d}_{+}$ and we have used $r_{0}=r_{\mathrm{m}}-5\times 10^{-5}M$ for numerical calculations of $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs.
+(29) and (30) as $r_{0}=r_{\mathrm{m}}+3\times 10^{-5}M$ for numerical calculations to get the values of $\bar{c}_{+}$ and $\bar{d}_{+}$ and we have used $r_{0}=r_{\mathrm{m}}-2.4\times 10^{-5}M$ for numerical calculations of $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs.
 (29) and (30) to get the values of $\bar{c}_{-}$ and $\bar{d}_{-}$.
-In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{c}_{+}\sim 6.01322$ in our numerical method and its difference from the value of $\bar{c}_{+}\sim 6.01316$ in Chiba and Kimura Chiba:2017nml is tiny but we have found that the value $\bar{c}_{+}\sim 4.95196$ obtained by the semi-analytic calculation in Tsukamoto:2020iez should be modified.
+In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{c}_{+}\sim 6.01324$ in our numerical method and its difference from the value of $\bar{c}_{+}\sim 6.01316$ in Chiba and Kimura Chiba:2017nml is tiny but we have found that the value $\bar{c}_{+}\sim 4.95196$ obtained by the semi-analytic calculation in Ref.
+Tsukamoto:2020iez should be modified.
 Thus, we conclude that our extent Eiroa, Romero, and Torres’s method to gravitational lensing of the rays just inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ gives the correct values of $\bar{c}_{\pm}$ and $\bar{d}_{\pm}$ and that the semi-analytic method considered by the author previously in Ref.
 Tsukamoto:2020iez gives the invalid values of $\bar{c}_{+}$ while it gives the correct values of $\bar{d}_{+}$.
-We comment on constraints on metrics from the shadow observations of supermassive black hole candidates at the centers of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp .
+We will revisit to modify the semi-analytic calculation to obtain $\bar{c}_{\pm}$ correctly in forthcoming work Igata:2026ivq .
+On this paper, we have concentrated on the extension of Eiroa, Romero, and Torres’s method to gravitational lensing of the rays inside and outside of the marginally unstable photon sphere.
+One may expect that we can apply it to the deflection angles (2) in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ in spacetimes with a photon sphere and an antiphoton sphere.
+As shown in Refs.
+Shaikh:2019itn ; Tsukamoto:2021fsz , however, $r_{0}$ does not approach to $r_{\mathrm{m}}$ but another value in the limit $b\rightarrow b_{\mathrm{m}}-0$.
+Thus, the strong deflection limit $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs.
+(27)-(30) is invalid and we need another extension of Eiroa, Romero, and Torres’s method for the deflection angles (2) in the strong deflection limit $b\rightarrow b_{\mathrm{m}}-0$ and it is left as a future work.
+VI Discussion
+The observed ring images of the supermassive black hole candidates at the centers of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp are consistent with lensing images formed by synchrotron radiations from a hot plasma near the black holes according to their ray-traced general-relativistic magnetohydrodynamic (GRMHD) simulations but the exotic compact objects such as naked singularity spacetimes, regular compact objects, and wormholes are not excluded Akiyama:2019cqa ; EventHorizonTelescope:2022wkp .
+This is because there are too many parameter regions for the exotic compact objects Akiyama:2019cqa ; EventHorizonTelescope:2022wkp and astrophisical models Gralla:2020pra to exclude the exotic cases.
+We comment on constraints on metrics from the shadow observations of supermassive black hole candidates at the centers of our galaxy and M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp .
 The Event Horizon Telescope collaborations EventHorizonTelescope:2021dqv ; EventHorizonTelescope:2022xqj constrained the parameter of the metrics of spacetimes from the observations and the sizes of the photon spheres under assumptions for simplify.
-Thus, the constraints.
-which depend on the assumptions.
-are not very reliable Gralla:2020pra ; Tsukamoto:2024gkz and the constraints on the metrics should be confirmed by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr .
+Thus, the constraints, which depend on the assumptions are not very reliable Gralla:2020pra ; Tsukamoto:2024gkz and the constraints on the metrics should be confirmed by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr .
+[…]
+A compact object with a marginary unstable photon sphere can be interesting exotic cases since it can be one of the most different cases from a Schwarzschild black hole among compact objects with diverging deflection angles.
+For the comparisons of the observable, we have to treat the rays inside of the marginally unstable photon sphere also since they can brighter than rays nearly outside the marginally unstable photon sphere.
+In Refs.
+Tsukamoto:2020iez and Chiba:2017nml , they do not give formulas for inside of the marginary unstable photon sphere while we and Ref.
+Sasaki:2025web do that.
+In Ref.
+Sasaki:2025web , we can see a complete and reliable treatment in the Reissner-Nordström spacetime case.
+Our approach in this paper will be gives larger errors than Ref.
+Sasaki:2025web but it has a merit that it is easy to apply to other spacetimes.
+The precise values in Ref Sasaki:2025web in the Reissner-Nordström spacetime case will be a milestone to improve our method more.
+We, however, emphasize that the larger errors in our numerical values in Table I does not affect our discussion as we seen below.
+[…]
+[…]
+Table 1: The coefficients $\bar{c}_{\pm}$ and the terms $\bar{d}_{\pm}$ in the deflection angles of the rays deflected inside and outside of the marginally unstable photon spheres in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ in the Reissner-Nordström (RN) spacetime obtained in our numerical method, by Sasaki Sasaki:2025web , and a semi-analytical method by Tsukamoto Tsukamoto:2020iez and the ones in the Hayward spacetime in our method, by Chiba and Kimura Chiba:2017nml , and in the semi-analytical method by Tsukamoto Tsukamoto:2020iez , and observable calculated by them are shown in the case with distances $d_{\mathrm{OS}}=16$kpc and $d_{\mathrm{OL}}=d_{\mathrm{LS}}=8$kpc, and the mass $M=4\times 10^{6}M_{\odot}$.
+The reduced critical impact parameter $b_{\mathrm{m}}/M$, the diameter of the marginally unstable photon sphere $2\theta_{\infty}$, the diameters of the Einstein rings $2\theta_{\mathrm{E}1\pm}$, $2\theta_{\mathrm{E}2\pm}$, and $2\theta_{\mathrm{E}3\pm}$ with the winding number $n=1$, 22, and $3$, respectively, the difference of the radii $\bar{s}_{\pm}=\theta_{\mathrm{E}1\pm}-\theta_{\infty}$, the total magnifications of the pair images $\mu_{1\mathrm{tot}\pm}(\phi)$, $\mu_{2\mathrm{tot}\pm}(\phi)$, and $\mu_{3\mathrm{tot}\pm}(\phi)$ with the source angle $\phi=1$ arcsecond, and with $n=1$, 22, and $3$, respectively, and the ratio of the magnifications $\bar{r}_{\pm}$ are shown.
+RN in our method RN in Sasaki:2025web RN in Tsukamoto:2020iez Hayward in our method Hayward in Chiba:2017nml Hayward in Tsukamoto:2020iez
+$b_{\mathrm{m}}/M$ $3\sqrt{6}/2$ $3\sqrt{6}/2$ $3\sqrt{6}/2$ $25\sqrt{5}/12$ $25\sqrt{5}/12$ $25\sqrt{5}/12$
+[…]
+$\bar{c}_{+}$ 6.67748 6.67748 5.49892 6.01324 6.01316 4.95196
+$\bar{d}_{+}$ -5.59011 -5.59108 -5.59108 -5.60958 - -5.62607
+[…]
+[…]
+[…]
+[…]
+$\mu_{1\mathrm{tot}+}(\phi)\times 10^{19}$ 1063.82 1063.19 324.595 888.554 - 271.347
+$\mu_{2\mathrm{tot}+}(\phi)\times 10^{19}$ 52.8675 52.8477 16.4542 44.9320 - 13.9129
+$\mu_{3\mathrm{tot}+}(\phi)\times 10^{19}$ 6.58948 6.58765 2.05397 5.61662 - 1.7433
+$\bar{r}_{+}$ 17.3389 17.3350 16.9947 17.0324 - 16.7917
+$\bar{c}_{-}$ 11.5658 11.5657 - 10.4154 - -
+$\bar{d}_{-}$ -5.64071 -5.59108 - -5.67374 - -
+[…]
+[…]
+[…]
+[…]
+$\mu_{1\mathrm{tot}-}(\phi)\times 10^{19}$ 4518.59 4066.21 - 12798.4 - -
+$\mu_{2\mathrm{tot}-}(\phi)\times 10^{19}$ 1304.72 1328.28 - 1141.19 - -
+$\mu_{3\mathrm{tot}-}(\phi)\times 10^{19}$ 173.347 175.793 - 148.000 - -
+$\bar{r}_{-}$ 2.95597 2.61464 - 9.60632 - -
+Acknowledgements
+The author would like to be grateful to T.
+Igata, T.
+Sasaki, H.
+Asada, and an anonymous referee for their useful and valuable comments.
+The author also would like to thank to M.
+Saijo for his comments in the very early stage of this work.